OpenStudy (anonymous):

three cards are drawn from an ordinary deck of 52 cards. each card is replaced in the deck before the next card is drawn. what is the probability that at least one of the cards was a spade

OpenStudy (anonymous):

Ok first how many spades are in a deck of cards

OpenStudy (anonymous):

13

OpenStudy (anonymous):

52C3

OpenStudy (anonymous):

soo divde 52 cards by 13 cards

OpenStudy (anonymous):

what do i put in the numerator?

OpenStudy (anonymous):

it says at least 1 spade so im confused

OpenStudy (anonymous):

round it to the nearest whole number

OpenStudy (anonymous):

P = 1 - P(no spade cards)

OpenStudy (anonymous):

im so confused now lol

OpenStudy (anonymous):

at least 1 spade card means P(1 spade) + P(2 spades) + P(3 spades). If you do this way, you'd have to calculate 3 times. But instead, just calculate the complement and subtract from 1

OpenStudy (anonymous):

wont i get a negative number if i subtract from 1?

OpenStudy (anonymous):

No, you won't

OpenStudy (anonymous):

what

OpenStudy (anonymous):

what would you get if you do P(no spade cards) ?

OpenStudy (anonymous):

39/52

OpenStudy (anonymous):

oh i have to subtract the fraction

OpenStudy (anonymous):

1-(39/52)=1/4

OpenStudy (anonymous):

you are drawing 3 cards, not 1 card

OpenStudy (anonymous):

1 - (39/52)^3